15 cm3 of 0.2 mole solution hydroxide react with 10cm3 of 0.15 mole of phosphoric acid H3PO4. Find the formula of the salt formed?
1
Expert's answer
2017-04-29T09:06:08-0400
Number of moles of acid and hydroxide equal:
n(H3PO4) = C·V = 0.15·0.01 = 0.0015 (moles)
n(Me(OH)x) = C·V = 0.2·0.015 = 0.003 (moles)
Thus, n(H3PO4): n(Me(OH)x) = 1:2
The balanced chemical equation is:
2 Me(OH)x + H3PO4 → Me2HPO4 + 2H2O
So the metal has a valence I, and general formula of hydroxide is MeOH.
So the formula of formed salt is Me2HPO4.

Answer: Me2HPO4.